# Jean DELHAYE

### v2

@@ 4. Restricting the study to a smaller algebra
Let us write
$\chi_{\mathbf{n}}^{\epsilon}=z^{[\epsilon_{0}]_{-}}\chi_{n_{1}}z^{\epsilon_{1}}...z^{\epsilon_{p-1}}\chi_{n_{p}}z^{[\epsilon_{p}]_{+}}.$
From there, we simply compute
$\sqrt{2}x\cdot\chi_{\mathbf{n}}^{\epsilon}$ $={(z^{[\epsilon]_{-}}\chi_{1}z^{[\epsilon]_{+}}+z^{[-\epsilon]_{-}}\chi_{1}z^{[-\epsilon]_{+}})z^{[\epsilon_{0}]_{-}}\chi_{n_{1}}z^{\epsilon_{1}}...z^{\epsilon_{p-1}}\chi_{n_{p}}z^{[\epsilon_{p}]_{+}}}$
$={z^{[\epsilon]_{-}}\chi_{1}z^{\epsilon}\chi_{n_{1}}z^{\epsilon_{1}}...z^{\epsilon_{p-1}}\chi_{n_{p}}z^{[\epsilon_{p}]_{+}}+z^{[-\epsilon]_{-}}\chi_{1}\chi_{n_{1}}z^{\epsilon_{1}}...z^{\epsilon_{p-1}}\chi_{n_{p}}z^{[\epsilon_{p}]_{+}}}$
$={\chi_{(1,\mathbf{n})}^{\epsilon}+z^{[-\epsilon]_{-}}(\chi_{n_{1}+1}+\chi_{n_{1}-1})z^{\epsilon_{1}}...z^{\epsilon_{p-1}}\chi_{n_{p}}z^{[\epsilon_{p}]_{+}}}$
$={\chi_{(1,\mathbf{n})}^{\epsilon}+\chi_{\mathbf{n}^{+}}^{-\epsilon}+\chi_{\mathbf{n}^{-}}^{-\epsilon}}.$
The following proposition shows that for any $m$, the element $x_{m}$ is a linear combination of irreducible characters whose tuples are compositions of $m$. This implies not only that all the information of the limit-process is contained within this algebra but it also gives the existence of a conditional expectation $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N}^{+})_{00}$ leaving both the Haar state and the limit-process invariant (seeing as the $x_{m}$’s constitute an orthonormal family for the Haar state).
Proposition 4.3.
The following assertions hold.
  1. (i)
    We have
    $\frac{x_{m}}{2^{m/2}}=\frac{1}{2^{m}}\sum_{\begin{subarray}{c}\mathbf{n}\in\Pi_{m},\\ \epsilon=\pm 1\end{subarray}}\chi^{\epsilon}_{\mathbf{n}},\quad m\in\mathbb{N}^{*}.$
  2. (ii)
    $\mathbb{F}[\chi_{\mathbf{n}}^{\epsilon}]=\frac{\xi_{|\mathbf{n}|}}{2^{|\mathbf{n}|/2}}.$
Proof.
(i) The result is clear for $m=1,2$, assume the result to hold up until some rank $m\geq 2$. To simplify computations, let us write $y_{\mathbf{n}}=\chi^{1}_{\mathbf{n}}+\chi^{-1}_{\mathbf{n}}$. Using Lemma 4.2 we have
$\frac{x_{m+1}}{2^{(m+1)/2}}$ $=\frac{x\cdot x_{m}-x_{m-1}}{2^{(m+1)/2}}$
$=\frac{x}{2^{1/2}\cdot 2^{m}}\sum_{\mathbf{n}\in\Pi_{m}}y_{\mathbf{n}}-\frac{1}{2\cdot 2^{m-1}}\sum_{\mathbf{n}\in\Pi_{m-1}}y_{\mathbf{n}}$
$=\frac{1}{2^{m+1}}\sum_{\mathbf{n}\in\Pi_{m}}y_{(1,\mathbf{n})}+y_{\mathbf{n}^{+}}+y_{\mathbf{n}^{-}}-\frac{1}{2^{m}}\sum_{\mathbf{n}\in\Pi_{m-1}}y_{\mathbf{n}}$
$=\frac{1}{2^{m+1}}\sum_{\mathbf{n}\in\Pi_{m+1}}y_{\mathbf{n}}+\frac{2}{2^{m+1}}\sum_{\mathbf{n}\in\Pi_{m-1}}y_{\mathbf{n}}-\frac{1}{2^{m}}\sum_{\mathbf{n}\in\Pi_{m-1}}y_{\mathbf{n}}$
$=\frac{1}{2^{m+1}}\sum_{\mathbf{n}\in\Pi_{m+1}}y_{\mathbf{n}}.$
This shows the result by induction.
(ii) Let $\mathrm{L}^{2}(U_{N}^{+})_{0}$ and $\mathrm{L}^{2}(U_{N}^{+})_{00}$ be the $\mathrm{L}^{2}$-spaces of $\mathcal{O}(U_{N}^{+})_{0}$ and $\mathcal{O}(U_{N}^{+})_{00}$ respectively and $\mathbb{F}^{\prime}$ be the orthogonal projection $\mathrm{L}^{2}(U_{N}^{+})_{0}\to\mathrm{L}^{2}(U_{N}^{+})_{00}$. Let us show that $\mathbb{F}^{\prime}$ satisfies the relation stated in the proposition, it suffices to show that
$h(\chi_{\mathbf{n}}^{\epsilon}\cdot x_{m})=\frac{\delta_{|\mathbf{n}|,m}}{2^{m/2}},\quad\mathbf{n}\in\bigsqcup_{p\geq 1}\mathbb{N}^{*p},m\in\mathbb{N}^{*}.$
An easy computation enables us to see that the latter relation holds whenever $m\leq 2$. Assume the relation to hold up to some rank $m\geq 2$. Using the traciality of the Haar state and the first item, we have
$h(\chi_{\mathbf{n}}^{\epsilon}\cdot x_{m+1})$ $=h(x\cdot\chi_{\mathbf{n}}^{\epsilon}\cdot x_{m})-h(\chi_{\mathbf{n}}^{\epsilon}\cdot x_{m-1})$
$=\frac{h\big{(}(\chi_{(1,\mathbf{n})}^{\epsilon}+\chi_{\mathbf{n}^{+}}^{-\epsilon}+\chi_{\mathbf{n}^{-}}^{-\epsilon})\cdot x_{m}\big{)}}{\sqrt{2}}-\frac{\delta_{|\mathbf{n}|,m-1}}{2^{(m-1)/2}}$
$=\frac{2\delta_{|\mathbf{n}|+1,m}+\delta_{|\mathbf{n}|-1,m}}{2^{(m+1)/2}}-\frac{\delta_{|\mathbf{n}|,m-1}}{2^{(m-1)/2}}$
$=\frac{\delta_{|\mathbf{n}|,m+1}}{2^{(m+1)/2}}.$
This proves the relation by induction on $m$. It follows from [1, Thm 9.1.2] that $\mathbb{F}=\mathbb{F}^{\prime}\circ\mathbb{E}$, where we have restricted $\mathbb{F}^{\prime}$ to $\mathcal{O}(U_{N}^{+})_{0}$, is the desired conditional expectation. ∎
Finally, let us end this subsection by mentioning that the closure of $\mathcal{O}(U_{N}^{+})_{00}$ in $C(U_{N}^{+})$ is isomorphic to $C([-N,N])$. Indeed, it suffices to see that the spectrum of $x/\sqrt{2}$ is $[-N,N]$. The inclusion $\sigma(x/\sqrt{2})\subset[-N,N]$ follows from the fact that $\|x/\sqrt{2}\|_{C(U_{N}^{+})}\leq N$. On the other hand, there exists a surjective C*-homomorphism $C(U_{N}^{+})\to C(O_{N}^{+}),u_{ij}\mapsto o_{ij}$. In particular, $x/\sqrt{2}$ is sent to $\chi_{1}$ which is well-known to have spectrum $[-N,N]$ (see [9, Lem 4.2]), hence the other inclusion.
5. The limit profile
$\psi_{t}(\chi_{2})=(N^{2}-1)e^{-2t(N\alpha-\beta)/(N^{2}-1)}\neq N^{2}e^{-2t(N\alpha+\beta)/N^{2}}=\psi_{t}(\chi_{(1,1)}).$
Let us first determine the limit profile of the process defined by $\widetilde{\psi}_{t}=\psi_{t}\circ\mathbb{F}$.
Theorem 5.1.
We have
$d_{\mathrm{TV}}\big{(}\widetilde{\psi}_{t_{c}},h\big{)}\underset{N\to\infty}{\longrightarrow}d_{\mathrm{T}V}(\eta_{c},\nu_{\mathrm{SC}}),\quad c>0$
and
$\underset{N\to\infty}{\lim\inf}\,d_{\mathrm{TV}}\big{(}\widetilde{\psi}_{t_{c}},h\big{)}\geq d_{\mathrm{T}V}(\eta_{c},\nu_{\mathrm{SC}}),\quad c<0,$
Proof.
Up to renormalising the process, we may assume that $\alpha=1$. Recall that the process being $\mathbb{F}$-invariant, we may restrict the study to the commutative algebra $\mathcal{O}(U_{N}^{+})_{00}$ for which the $x_{m}$’s form an orthonormal basis.
The proof of the first convergence closely follows the reasoning in the proof [23, Prop 3.10], with a few modifications. Let us outline the elements that differ. Let $c>0$. It follows from Proposition 3.8 that $\widetilde{\psi}_{t_{c}}$ has an $\mathrm{L}^{1}$-density (at least asymptotically defined) given by
$\widetilde{f}_{t_{c}}=\sum_{m\geq 0}{\widetilde{\psi}_{t_{c}}(x_{m})}x_{m}=\sum_{m\geq 0}\frac{1}{2^{m/2}}\left(\sum_{\begin{subarray}{c}\mathbf{n}\in\Pi_{m},\\ \epsilon=\pm 1\end{subarray}}\psi_{t_{c}}(\chi^{\epsilon}_{\mathbf{n}})\right)x_{m}.$
Using the isomorphism $\mathcal{O}(U_{N}^{+})_{00}\to\mathbb{C}[X],x_{m}\mapsto P_{m}$ sending the Haar state to the semi-circle distribution, we have
$d_{\mathrm{TV}}(\widetilde{\psi}_{t_{c}},h)=\frac{1}{2}\|\widetilde{f}_{t_{c}}-1\|_{1,h}=\frac{1}{2}\left\|\sum_{m\geq 1}\widetilde{\psi}_{t_{c}}(x_{m})P_{m}\right\|_{1,\nu_{\mathrm{SC}}}.$
In parallel, note that for any tuple $\mathbf{n}=(n_{1},...,n_{p})$, we have $d_{\mathbf{n}}e^{-t_{c}\lambda_{\mathbf{n}}}\to e^{-\widetilde{c}|\mathbf{n}|}$ as $N\to\infty$ where $\widetilde{c}=c+\ln\sqrt{2}$ (this follows from Lemma 4.1 and Proposition 4.3). This further implies that for any given $m\in\mathbb{N}$, we have
$\psi_{t_{c}}(x_{m})$ $=\frac{1}{2^{m/2}}\sum_{\begin{subarray}{c}\mathbf{n}\in\Pi_{m},\\ \epsilon=\pm 1\end{subarray}}\psi_{t_{c}}(\chi_{\mathbf{n}}^{\epsilon})=\frac{1}{2^{m/2}}\sum_{\begin{subarray}{c}\mathbf{n}\in\Pi_{m},\\ \epsilon=\pm 1\end{subarray}}d_{\mathbf{n}}e^{-t_{c}|\mathbf{n}|}\exp\left(t_{c}\beta\left(\frac{\lambda_{\mathbf{n}}}{N}-\frac{2\mathfrak{e}_{\mathbf{n}}+\mathfrak{p}_{\mathbf{n}}}{N^{2}}\right)\right)$
$=\frac{1}{2^{m/2}}\sum_{\begin{subarray}{c}\mathbf{n}\in\Pi_{m},\\ \epsilon=\pm 1\end{subarray}}d_{\mathbf{n}}e^{-t_{c}|\mathbf{n}|}\exp\left((N\ln(\sqrt{2}N)+cN)O(N^{-3})\right)=\frac{1}{2^{m/2}}\sum_{\begin{subarray}{c}\mathbf{n}\in\Pi_{m},\\ \epsilon=\pm 1\end{subarray}}d_{\mathbf{n}}e^{-t_{c}m+o(1)}$
$\underset{N\to\infty}{\longrightarrow}2^{m/2}e^{-\widetilde{c}m}=e^{-cm}.$
On the other hand, for any fixed $m\in\mathbb{N}$, we have
$\|\psi_{t_{c}}(x_{m})x_{m}\|_{1}$ $\leq\|\psi_{t_{c}}(x_{m})x_{m}\|_{2}$
$=\frac{1}{2^{m/2}}\sum_{\begin{subarray}{c}\mathbf{n}\in\Pi_{m},\\ \epsilon=\pm 1\end{subarray}}\psi_{s_{t}}(\chi^{\epsilon}_{\mathbf{n}})$
$\leq\frac{1}{2^{m/2}}\sum_{\begin{subarray}{c}\mathbf{n}\in\Pi_{m},\\ \epsilon=\pm 1\end{subarray}}d_{\mathbf{n}}e^{-t_{c}(1-\beta/N)\lambda_{\mathbf{n}}}$
$\leq 2^{m/2}\frac{q^{-m}}{(1-q^{2})^{m}}e^{-t_{c}(1-\beta/N)m/N}.$
As mentioned in the proof of Proposition 3.8, the latter ratio can uniformly in $N$ (at least from a certain rank) be bounded by a constant strictly less than one. This implies that we may exchange the sum over $m$ and the limit in $N$, yielding
$\lim_{N\to\infty}d_{\mathrm{TV}}(\widetilde{\psi}_{t_{c}},h)=\frac{1}{2}\left\|\sum_{m\geq 1}e^{-cm}P_{m}\right\|_{1,\nu_{\mathrm{sc}}}=d_{\mathrm{TV}}(\eta_{c},\nu_{\mathrm{SC}}).$
Now, for the second result, fix $c<0$ and consider the states $\widetilde{\psi}_{t_{c}}$ and $h$ in their measure form as $\mu_{t_{c}}$ and $\nu_{\mathrm{SC}}$. We have already established that $\mu_{t_{c}}$ converges to $\eta_{c}$ in moments as $N\to\infty$. Taking $N$ large enough so that $t_{c}\beta\leq-cN^{2}$, using the estimates from [17, Lem 1.7], Lemma 3.7 and the fact that $d_{n}\leq N^{n}$, we have for any $m\in\mathbb{N}$
$\mu_{t_{c}}(P_{m})\leq 2^{m/2}N^{m}e^{-t_{c}(1-\beta/N)m/N}\leq e^{-2cm}=\eta_{2c}(P_{m}).$
In particular, this implies that:
$\mu_{t_{c}}(X^{2n})\leq\eta_{2c}(X^{2n}),\quad\text{for all }n\in\mathbb{N}.$
This follows from the fact that $X^{2n}$ can be expressed as a linear combination of $P_{m}$’s with positive coefficients. Since $\eta_{2c}$ is supported on the interval $[-\gamma_{2c},\gamma_{2c}]$, where $\gamma_{2c}=e^{2c}+e^{-2c}$, it follows that $\mu_{t_{c}}$ must also be supported on this interval (as the support of a measure can be estimated by bounding its moments; see, for example, the beginning of [32, Lect 2]). Given that moment convergence with compact support implies weak convergence, we have:
$\mu_{t_{c}}(B)\xrightarrow[N\to\infty]{}\eta_{c}(B),$
for any continuity set $B$ of $\eta_{c}$, i.e., a set such that $\eta_{c}(\partial B)=0$. The final step is to identify such a set that maximizes the total variation distance between $\eta_{c}$ and $\nu_{\mathrm{SC}}$. It suffices to take:
$B_{c}=\{x\in[-2,2]:g_{c}(x)\leq 1\},$
where $g_{c}$ is the density of the absolutely continuous part of $\eta_{c}$ with respect to the semicircular distribution. With this choice of $B_{c}$, we now conclude the argument.
$d_{\mathrm{TV}}(\mu_{t_{c}},\nu_{\mathrm{SC}})\geq\nu_{\mathrm{SC}}(B_{c})-\mu_{t_{c}}(B_{c})\underset{N\to\infty}{\longrightarrow}\nu_{\mathrm{SC}}(B_{c})-\eta_{c}(B_{c})=d_{\mathrm{TV}}(\eta_{c},\nu_{\mathrm{SC}}).$
We can now conclude for the pure Brownian motion.
Corollary 5.2.
We have
$d_{\mathrm{TV}}\big{(}\psi_{t_{c}},h\big{)}\underset{N\to\infty}{\longrightarrow}d_{\mathrm{T}V}(\eta_{c},\nu_{\mathrm{SC}}),\quad c>0$
and
$\underset{N\to\infty}{\lim\inf}\,d_{\mathrm{TV}}\big{(}\psi_{t_{c}},h\big{)}\geq d_{\mathrm{T}V}(\eta_{c},\nu_{\mathrm{SC}}),\quad c<0,$
where $t_{c}=N\ln(\sqrt{2}N)/\alpha+cN/\alpha$. In particular, the Brownian motion has cutoff at time $t_{N}=N\ln(\sqrt{2}N)/\alpha$.
Proof.
Fix $c>0$, for the first convergence it suffices to see that $d_{\mathrm{TV}}\big{(}\psi_{t_{c}},\widetilde{\psi}_{t_{c}}\big{)}\to 0$ as $N\to\infty$. This is obvious as
$d_{\mathrm{TV}}\big{(}\psi_{t_{c}},\widetilde{\psi}_{t_{c}}\big{)}\leq\sum_{\epsilon,\mathbf{n}}|\psi_{t_{c}}(\chi_{\mathbf{n}}^{\epsilon})-\widetilde{\psi}_{t_{c}}(\chi_{\mathbf{n}}^{\epsilon})|^{2}\leq\|f_{t_{c}}\|_{2}^{2}+\|\widetilde{f}_{t_{c}}\|_{2}^{2}\leq 2\sum_{m\geq 1}e^{-cm}$
where the last inequality holds for $N$ large enough (similarly to the proof of Proposition 3.8). By dominated convergence, we may exchange summation and limit in $N$ yielding the convergence.
For the second one, simply note that for any $c\in\mathbb{R}$, we have
$d_{\mathrm{TV}}(\psi_{t_{c}},h)=\frac{1}{2}\|\psi_{t_{c}}-h\|_{FS}\geq\frac{1}{2}\|\mathbb{F}\circ\psi_{t_{c}}-\mathbb{F}\circ h\|_{FS}=d_{\mathrm{TV}}(\widetilde{\psi}_{t_{c}},h).$
Let us conclude this section by discussing the result. The limit profile on the right ($c>0$) is straightforward to compute due to the inherent absolute continuity in this region. However, outside of this region, absolute continuity is lost. It is worth recalling that such a loss of absolute continuity was already observed in [23], where all processes studied exhibited an atom as the singular part when absolute continuity was no longer present. P. Biane, in [6, Sec 12.2], introduced an interesting method for recovering the absolutely continuous part of a measure when only its moments are known, provided that one can identify its atom. In our case, the singular part is more complex than a single atom, and it remains unclear whether Biane’s method could be applied here. We conjecture, however, that the singular part lies outside the support of $\nu_{\mathrm{SC}}$, and that there is true $\mathrm{L}^{1}$-convergence. Specifically, we believe that even for $c<0$, the absolutely continuous part of the process converges to that of $\eta_{c}$ in $\mathrm{L}^{1}(\nu_{\mathrm{SC}})$.
6. Non-uniqueness of limit profiles
When examining a limit profile for a family of (quantum) groups, each paired with a central Lévy process $(G_{N},\varphi^{N})_{N\in\mathbb{N}}$, as defined in Definition 2.7, it is pertinent to consider how the profile depends on the structure of $G_{N}$ and the dynamics of $\varphi^{N}$. Although the impact of the group’s representation theory on the profile is obvious (see the profiles studied in [23, 30, 33] and this paper), the degree to which the process itself influences the limit profile is not clear. This raises a question: is the temporal aspect the only significant parameter through which the process affects the cutoff behavior?
This section presents an example to demonstrate that different central Lévy processes on the same quantum group can lead to distinct limit profiles. Consider the orthogonal quantum group $O_{N}^{+}$ and its associated Brownian motion, which typically has a limit profile function $c\mapsto d_{\mathrm{TV}}(\eta_{c},\nu_{\mathrm{SC}})$. We define a new Lévy process on $O_{N}^{+}$, designed to exhibit a different limit profile.
The central generating functional for this new process is defined by:
$L:\chi_{n}\mapsto-P^{\prime}_{n}(N)-\frac{P_{n}(N)-P_{n}(-N)}{2}=\left\{\begin{array}[]{ll}-P_{n}^{\prime}(N)&\text{if }n\text{ is even},\\ -P_{n}^{\prime}(N)-P_{n}(N)&\text{otherwise}.\end{array}\right.$
This is the functional with parameters $b=1$ and $\nu=N\delta_{-N}$, when considering the decomposition in Equation (3.1). The resulting Lévy process $\varphi_{t}$ is then given by:
$\varphi_{t}:\chi_{n}\longmapsto d_{n}e^{-t\lambda_{n}}c_{t,n}\quad\text{where }c_{t,n}=\left\{\begin{array}[]{ll}1&\text{if }n\text{ is even},\\ e^{-t}&\text{otherwise}.\end{array}\right.$
Theorem 6.1.
Fix $c\in\mathbb{R}$ and set $t_{N}:=N\ln(N)+cN$, we have
$d_{\mathrm{TV}}(\varphi_{t_{N}},h)\underset{N\to\infty}{\longrightarrow}d_{\mathrm{TV}}(\hat{\eta}_{c},\nu_{\mathrm{SC}}),$
where $\hat{\eta}_{c}$ is the unique distribution satisfying $\hat{\eta}_{c}(P_{n})=e^{-nc}$ if $n$ is even and 00 otherwise. The distribution $\hat{\eta}_{c}$ can be described by:
$\hat{\eta}_{c}=\mathbf{1}_{c\leq 0}(1-e^{2c})\frac{\delta_{\gamma_{c}}+\delta_{-\gamma_{c}}}{2}+\frac{e^{c}\gamma_{c}}{\gamma_{c}^{2}-x^{2}}\mathrm{d}\nu_{\mathrm{SC}}(x),\quad\text{with }\gamma_{c}=e^{c}+e^{-c}.$
Proof.
Fix $c\in\mathbb{R}$. Let us first prove that the distribution $\hat{\eta}_{c}$ as described in the theorem has the desired moments. If $c>0$, then clearly the series $\sum_{n\geq 0}e^{-2cn}P_{2n}$ belongs to $\mathrm{L}^{1}(\nu_{\mathrm{SC}})$. Applying twice the recursive formula satisfied by the Chebyshev polynomials confirms that the series is the density of $\hat{\eta}_{c}$ w.r.t $\nu_{\mathrm{SC}}$. If $c=0$, then we conclude by dominated convergence
$\hat{\eta}_{0}(P_{n})=\int_{-2}^{2}\frac{2P_{n}(x)}{2-x^{2}}\mathrm{d}\nu_{\mathrm{SC}}(x)=\lim_{c\to 0^{+}}\hat{\eta}_{c}(P_{n})=\left\{\begin{array}[]{cl}1&\hbox{if }n\hbox{ is even}\\ 0&\hbox{else},\end{array}\right.\quad n\in\mathbb{N}.$
Finally, if $c<0$, using the positive case, we have whenever $n$ is even
$\hat{\eta}_{c}(P_{n})$ $=(1-e^{2c})\frac{P_{n}(\gamma_{c})+P_{n}(-\gamma_{c})}{2}+e^{2c}\hat{\eta}_{-c}(P_{n})$
$=(1-e^{2c})\frac{e^{-(n+1)c}-e^{(n+1)c}}{e^{-c}-e^{c}}+e^{(n+2)c}=e^{-nc}.$
Clearly, the computation leads to zero when $n$ is odd.
Now to prove the limit profile recall that we can see $\varphi_{t_{N}}$ and $h$ as their measure counterparts $\mu_{t_{N}}$ and $\nu_{\mathrm{SC}}$ via the isomorphism $\chi_{n}\mapsto P_{n}$. It remains to see that the absolutely continuous part of $\mu_{t_{N}}$ w.r.t $\nu_{\mathrm{SC}}$ tends to that of $\hat{\eta}_{c}$ as $N\to\infty$ in total variation distance which can be done in a similar fashion as in the proof of [23, Thm 3.9]. ∎
Observe that the functions $f_{1}:c\mapsto d_{\mathrm{TV}}(\eta_{c},\nu_{\mathrm{SC}})$ and $f_{2}:c\mapsto d_{\mathrm{TV}}(\hat{\eta}_{c},\nu_{\mathrm{SC}})$ can be explicitly computed through integration, resulting in the following expressions:
$f_{1}(c)=\begin{cases}\frac{2(1-e^{2c})}{\pi}\arcsin\left(\frac{e^{-c}}{2}\right)+\frac{e^{-2c}+2}{4\pi}\sqrt{4e^{2c}-1},&\text{if }c\geq-\ln 2,\\ 1-e^{2c},&\text{if }c<-\ln 2,\end{cases}$
and
$f_{2}(c)=\begin{cases}\frac{\sqrt{1+e^{-2c}}\sqrt{3-e^{-2c}}}{2\pi}+\frac{e^{2c}-1}{\pi}\left(\arctan\left(\frac{\sqrt{1+e^{-2c}}(e^{c}-e^{-c})}{(e^{c}+e^{-c})\sqrt{3-e^{-2c}}}\right)-\arcsin\left(\frac{\sqrt{1+e^{-2c}}}{2}\right)\right),&\text{if }c\geq-\frac{\ln 3}{2},\\ 1-e^{2c},&\text{if }c<-\frac{\ln 3}{2}.\end{cases}$
Figure 1. Graph comparing the functions $f_{1}$ (red) and $f_{2}$ (blue).
They agree on infinitely many points but are indeed distinct. This implies that they are not equivalent modulo affine transformation. It follows from Remark 2.8 that the profile we exhibited is different.
Interestingly, although we might have anticipated that the new process would result in a ’faster shuffle’, in the sense that $f_{2}\leq f_{1}$, the two profiles initially overlap for a significant range before the impact of the ’faster shuffle’ becomes apparent.
Appendix
In this appendix, we give a description of Gaussian centralized generating functionals on $O_{N}^{+}$ and $U_{N}^{+}$. We will therefore focus our study on the latter quantum groups, we refer the reader to [19, Sec 1.5] for definitions in greater generality.
We would like to acknowledge that this result was previously achieved by different means by Franz, Freslon, and Skalski in an unpublished work providing a comprehensive description of Gaussian generating functionals on certain quantum groups.
@@ 6. Non-uniqueness of limit profiles
    $\eta(ab)=\rho(a)\eta(b)+\eta(b)\varepsilon(b),\quad a,b\in\mathcal{O}(\mathbb{G});$
@@
    $L(ab)=\varepsilon(a)L(b)+L(a)\varepsilon(b)+\langle\eta(a^{*}),\eta(b)\rangle,\quad a,b\in\mathcal{O}(\mathbb{G}).$
@@
$\eta_{2}(a)=U\eta_{1}(a)\rho_{2}(a)U=U\rho_{1}(a),\quad a\in\mathcal{O}(\mathbb{G}).$
@@
    $L(abc)=L(ab)\varepsilon(c)+L(ac)\varepsilon(b)+L(bc)\varepsilon(a)-\varepsilon(ab)L(c)-\varepsilon(ac)L(b)-\varepsilon(bc)L(a)$ for all $a,b,c\in\mathcal{O}(\mathbb{G})$;
@@
    $\eta(ab)=\varepsilon(a)\eta(b)+\eta(a)\varepsilon(b)$ for all $a,b\in\mathcal{O}(\mathbb{G})$.

### v3

@@ 7.2. Computation of the limit profile
Fix $r\in[0,\infty]$. First recall (see Proposition 5.1) that we may write
$\eta_{c}^{r}=f_{c}^{r}\mathrm{d}\nu_{\mathrm{SC}}+\eta_{c,\mathrm{sing}}$
with $\eta_{c,\mathrm{sing}}^{r}\perp\nu_{\mathrm{SC}}$ and
$f_{c}^{r}(x)=\mathbb{E}\left[\frac{1}{e^{-2c}R_{r}^{2}-xe^{-c}R_{r}+1}\right]\eta_{c,\mathrm{sing}}^{r}(A)=\mathbb{E}\left[\left(1-e^{2c}R_{r}^{-2}\right)_{+}\mathbf{1}_{A}(e^{-c}R_{r}+e^{c}R_{r}^{-1})\right],$
where $R_{r}=\cos(T_{r})$ for some random variable of Gaussian distribution $T_{r}\sim\mathcal{N}(0,2r)$ (with $T_{0}=0$ and $T_{\infty}\sim\mathrm{Unif}([0,2\pi])$). Since $f_{c}^{r}\downarrow 0$ uniformly as $c\to$-$\infty$, the total variation distance between $\eta_{c}^{r}$ and $\nu_{\mathrm{SC}}$ is carried out by the weight of the singular part for $c$ close enough to -$\infty$. In other words, $g_{r}$ coincides with $c\mapsto\eta_{c,\mathrm{sing}}^{r}(\mathbb{R})$ in an interval of the form $(-\infty,c_{r})$.
Fix $r,s>0$ and assume that $g_{r}$ and $g_{s}$ are equivalent modulo affine transformation, i.e.
$g_{r}(ac+b)=g_{s}(c),\quad c\in\mathbb{R},$
for some $a>0$ and $b\in\mathbb{R}$. Following Lemma 7.4, this implies
$1+A_{r}e^{ac+b}+B_{r}e^{2(ac+b)}+O(e^{3(ac+b)})=1+A_{s}e^{c}+B_{s}e^{2c}+O(e^{3c}),\quad c\to\infty.$
This already implies $a=1$, we are left with the system
$\left\{\begin{array}[]{cc}A_{r}e^{b}&=A_{s}\\ B_{r}e^{2b}&=B_{s}\end{array}\right$.
This implies $A_{r}^{2}/B_{r}=A_{s}^{2}/B_{s}$ which forces $r=s$ as the function
$x\longmapsto\frac{A_{x}^{2}}{B_{x}}=\frac{4}{\pi^{2}}(e^{2x^{2}}-e^{-2x^{2}})^{2},$
is injective on $\mathbb{R}_{+}$.
Assume $g_{0}$ and $g_{r}$ (for $r\in(0,\infty)$) are equivalent modulo affine transformation, this would imply (using again Lemma 7.4)
$1-e^{2(ac+b)}=1+A_{s}e^{c}+B_{s}e^{2c}+O(e^{3c}),\quad c\to-\infty.$
for some $a>0$ and $b\in\mathbb{R}$, which is impossible. So $g_{0}$ is not equivalent modulo affine transformation to $g_{r}$ for $r\neq 0$, similarly it is not equivalent modulo affine transformation to $g_{\infty}$.
Assume now that $g_{r}$ and $g_{\infty}$ are equivalent modulo affine transformation for some $r\in(0,\infty)$. This implies, using Lemma 7.4:
$1-\frac{4e^{c}}{\pi}+\frac{e^{3c}}{3\pi}+O(e^{5c})=1+A_{r}e^{ac+b}+B_{r}e^{2(ac+b)}+O(e^{3(ac+b)}),\quad c\to-\infty.$
for some $a>0$ and $b\in\mathbb{R}$. The second order implies $a=1$ and the third $a=3/2$ which is a contradiction. ∎
Below is a plot of the different cutoff profiles $g_{r}$ ($0\leq r\leq\infty$).
Figure 1. Total variation distance $g_{r}$ for increasing $r$ (top to bottom).
Let us conclude this section by discussing the result of Theorem 7.3. The limit profile on the right ($c>0$) is straightforward to compute thanks to the absolute continuity in this region. However, outside of this region, absolute continuity is lost. It is worth recalling that such a loss of absolute continuity was already observed in [FTW21], where all processes studied exhibited an atom as the singular part when absolute continuity was lost. P. Biane, in [Bia08, Sec 12.2], introduced an interesting method for recovering the absolutely continuous part of a measure when only its moments are known, provided that one can identify its atom. In our case, the singular part is more complex than a single atom, and it remains unclear whether Biane’s method could be applied here. We conjecture, however, that the singular part lies outside the support of $\nu_{\mathrm{SC}}$, and that there is true $\mathrm{L}^{1}$-convergence. Specifically, we believe that even for $c<0$, the absolutely continuous part of the process converges to that of $\eta_{c}$ in $\mathrm{L}^{1}(\nu_{\mathrm{SC}})$. Nonetheless, we have reasons to believe that the density of the absolutely continuous part (when $c<0$) is in $\mathrm{L}^{1}(\nu_{\mathrm{SC}})$ (by definition) but not in $\mathrm{L}^{2}(\nu_{\mathrm{SC}})$, which renders Biane’s method inapplicable here.
Appendix A
In this appendix, we give a description of Gaussian centralized generating functionals on $O_{N}^{+}$ and $U_{N}^{+}$. We will therefore focus our study on the latter quantum groups, we refer the reader to [FS16, Sec 1.5] for definitions in greater generality.
We would like to acknowledge that this result was previously achieved by different means by Franz, Freslon, and Skalski in [FFS24] (unavailable at the time of the first version of this work’s preprint) in which the authors provide a comprehensive description of Gaussian generating functionals on $O_{N}^{+}$ and $U_{N}^{+}$.
@@
    $\eta(xy)=\rho(x)\eta(y)+\eta(x)\varepsilon(y),\quad x,y\in\mathcal{O}(\mathbb{G});$
@@
    $L(xy)=\varepsilon(x)L(y)+L(x)\varepsilon(y)+\langle\eta(x^{*}),\eta(y)\rangle,\quad x,y\in\mathcal{O}(\mathbb{G}).$
@@
$\eta_{2}(x)=U\eta_{1}(x)\rho_{2}(x)U=U\rho_{1}(x),\quad x\in\mathcal{O}(\mathbb{G}).$
@@
    $L(xyz)=L(xy)\varepsilon(z)+L(xz)\varepsilon(y)+L(yz)\varepsilon(x)-\varepsilon(xy)L(z)-\varepsilon(xz)L(y)-\varepsilon(yz)L(x)$ for all $x,y,z\in\mathcal{O}(\mathbb{G})$;
@@
    $\eta(xy)=\varepsilon(x)\eta(y)+\eta(x)\varepsilon(y)$ for all $x,y\in\mathcal{O}(\mathbb{G})$.